OpenStudy (anonymous):

Find the exact area of the circle with the given radius. r=11.64m A.11.64m2 B.23.28m2 C.33.8724m2 D.135.4896m2

OpenStudy (ikram002p):

area of circle A radius. r is given by \(A=r^2\pi \) s.t \(\pi =3.14\)

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

135.4896m2 was right

OpenStudy (ikram002p):

^^ np